# “Haet Bhasha aur Diskrimineshun”: Phonetic Perturbations in Code-Mixed Hinglish to Red-Team LLMs

### v3

@@ 5 Results & Observations
We now present the results from our red-teaming experiments for all research questions described previously.
5.1 Success of Red-teaming Approach (RQ.1)
@@
In Table 1 we report the AASR for all prompt sets, models and jailbreak templates.
ChatGPT and Llama: The models are fairly robust to attacks in English, with AASR decreasing further when combined with the jailbreak templates. For ‘None’, AASR significantly increases in both the English $\rightarrow$ CM and CM $\rightarrow$ CMP transitions. Combining the CM or CMP prompts with the jailbreak templates again results in $\simeq$ 0 AASR in most cases, further reinforcing their robustness to template bases attacks. We note that Sandbox is the best-performing jailbreak template for ChatGPT, while AIM is the best for Llama.
Gemma and Mistral: Both models report very high AASR across the templates and prompt sets. For ‘None’ on the English set, both models already yield high AASR, showing their vulnerability to classic template-based attacks despite evidence of harm. With the CM set for ‘None’, Gemma becomes highly complicit, while Mistral also shows a significant AASR jump. When combined with the templates, models reach up to 0.99 AASR for both the English and the CM sets. In the CM $\rightarrow$ CMP transition, AASR stays nearly the same across the templates, with a slight drop in the ‘None’ case.
Comparing the responses of the human annotators with the LLM judge, we obtain an ICC value of 0.87, indicating a high degree of agreement Bartko (1966).
Significance Testing: Next, we evaluate the benefits of CM over English and CMP over CM using the Wilcoxon test Wilcoxon (1992) for each model and jailbreak template individually($p$-value = 0.050.053. The findings are as follows.
  • ChatGPT: CM is beneficial for the None, OM, Sandbox and AIM configurations. CMP is beneficial only in the None case.
  • Llama: CM is beneficial for the None and AIM configurations, while CMP is beneficial only in the None case.
  • Gemma: CM is beneficial for the None and AIM configurations, while CMP does not provide statistically significant benefits in any configuration.
  • Mistral: CM is beneficial for the None case while CMP does not provide statistically significant benefits in any configuration.
Thus, we conclude that two out of the four models are successfully red-teamed using our input format and jailbreak template. Next, we attempt to understand how relevant the outputs are for each input prompt.
5.2 Relevance of Generated Outputs (RQ.2)
@@
ChatGPT and Llama: Both models maintain very high AARR across all jailbreak templates, with AARR $\simeq$ 1 in many cases. Irrespective of the template, both models report the AARR scores in the CM and CMP templates comparable to English in most cases, demonstrating their ability to correctly interpret code-mixed inputs even with misspellings. Thus, the majority of the responses are highly relevant to the prompt and, therefore, harmful.
Gemma and Mistral: Both models have generally high AARR scores on the English set across all templates. AARR drops significantly in the English $\rightarrow$ CM and CM $\rightarrow$ CMP transitions for Gemma and the CM $\rightarrow$ CMP transition for Mistral in most cases. Simultaneously, the AASR has a non-increasing trend except for ‘None’, demonstrating that both models undergo only minor AASR drops despite significant AARR drops in the prompt-set transitions. Thus, we conclude that while ChatGPT and Llama respond contextually to the harmful prompts from all sets, the language understanding of Gemma and Mistral degrades significantly in the CMP set, hence the low AARR. Despite this, both maintain high AASR due to their general vulnerability to classic template-based attacks.
5.3 Explaining Phonetic Perturbations (RQ.3)
@@
In Figures 24, we plot the sequence attribution scores using the method of Integrated Gradients Sundararajan et al. (2017), for a given input prompt in the English set, as well as its CM and CMP variants for the Llama-3-8B-Instruct LLM. We study the attribution scores for generating safe outputs– a higher token score implies higher contribution towards triggering the safety filter and vice versa. In Figure 2, we observe how the tokens of the prompt 3“hate”, “speech” and “discrimination” have a high attribution score at the embedding layer, 1st decoder layer and the 8th decoder layer. This means that these tokens are primarily responsible for generating the safe response from the LLM. “hate” and “speech” retain their high importance even at the 16th decoder layer. A similar observation can be made in Figure 3, for the same prompt written in a code-mixed (HI-EN) manner, where “hate” and “speech” are written in the English language. Thus, standard code-mixing may not be enough to bypass the safety filters of LLMs. On the other hand, in Figure 4, we note that the CMP variant generates radically different tokens– “hate” $\rightarrow$ “haet” tokenized as “ha” + “et” and, “discrimination” $\rightarrow$ “bhed bhav” tokenized as “b” + “hed” + “b” + “ha” + “av”. This leads to a change in the attribution scores as well, which are now low for the sensitive words and thus do not trigger the safety filters. A similar observation can be made for other prompts. Thus, based on our observations, we can state that phonetic perturbations lead to input tokenisation in a way that impacts the safety filters of LLMs, thus allowing attackers to generate harmful outputs.
5.4 Red-teaming multimodal models (RQ.4)
@@
We now discuss the results of our multimodal experiment for the image generation task in Table 2. We present example generations from ChatGPT for each category of input prompt in Figure 5.
ChatGPT: The model is quite robust to English attacks in the ‘Base’ (equivalent to ‘None’ in Table 1) case. AASR noticeably increases with CM, with a significant boost with CMP for both templates with consistently high AARR– a trend similar to the text generation task. VisLM outperforms ‘Base’ for all prompt sets, contrasting the text generation task in which jailbreak templates degrade AASR.
Gemini: For ‘Base’, AASR drops considerably in the English $\rightarrow$ CM transition, followed by a slight recovery with the CMP set. Conversely, there is a consistent, albeit minor increase in the transitions with VisLM, finally achieving the highest AASR on the CMP set. Similar to ChatGPT, we observe consistently high AARR for all configurations. This confirms the effectiveness of our approach even in multimodal settings.
Metric
Jailbreak
Template
ChatGPT Gemini
English CM CMP English CM CMP
AASR Base 0.20 0.29 0.65 0.3 0.19 0.25
VisLM 0.35 0.45 0.78 0.38 0.4 0.43
AARR Base 0.93 0.98 0.95 0.97 0.96 0.94
VisLM 1.00 0.98 0.94 0.98 0.98 0.96
Table 2: AASR and AARR scores for ChatGPT and Gemini across Base and VisLM jailbreak templates and all input sets, i.e., Standard English prompts, CM prompts, CMP prompts. Metric-wise maximum values in each column are in bold.
Jailbreak Template Category ChatGPT Gemini
@@
Base Religious Hate 0.10 0.25 0.75 0.4 0.15 0.3
Gore 0.45 0.70 0.85 0.1 0.15 0.15
Self-Harm 0.15 0.20 0.85 0.4 0.25 0.2
Casteist Hate 0.10 0.10 0.30 0.25 0.1 0.15
SM Toxicity 0.30 0.35 0.60 0.45 0.4 0.5
VisLM Religious Hate 0.30 0.40 0.90 0.55 0.5 0.65
Gore 0.75 0.70 0.90 0.05 0.05 0.15
Self-Harm 0.20 0.45 0.95 0.5 0.5 0.45
Casteist Hate 0.15 0.20 0.25 0.3 0.45 0.45
SM Toxicity 0.50 0.60 0.80 0.65 0.65 0.6
@@
Table 3: Category-wise AASR for the image generation task for ChatGPT and Gemini across both jailbreak templates and all input sets, i.e., Standard English prompts, CM prompts and CMP prompts. Template-wise maximum values for each column are in bold.
In Table 3, we present the AASR per input prompt category for the image generation task. For ‘Base’ and across all prompt sets, ChatGPT has the highest AASR for gore-related images and Gemini for social media toxicity. VisLM maintains this trend except for the CMP case in which the categories switch. Lastly, VisLM generally improves AASR over ‘Base’.
6 Discussion
In this study, we develop input prompts involving code-mixing and phonetic perturbations to red-team multimodal generative AI models, viz. LLMs and MLLMs. To aid our red-teaming efforts in LLMs, we propose Sandbox– extending Opposite Mode to simulate a resilience testing environment and a new template for MLLMs– VisGPT.
Our results from the LLM red-teaming effort reveal the impact of textese style attacks in probing the safety filters of LLMs. ChatGPT and Llama show resistance to standard jailbreak templates, exhibiting alignment against such attacks. However, this alignment is rendered inadequate when exposed to our novel technique. Both CM and CMP provide significant improvement in jailbreak success as compared to only English prompts. The experiments reveal that CMP substantially improves AASR in Table 1 for a vanilla setup without applying additional jailbreak templates, exposing how the safety alignment of models is almost hard-coded to certain templates. Additionally, we observe this resistance to jailbreak templates breaking with the introduction of previously unseen templates, as seen in the substantial success of VisLM over Base in the image generation task. On the other hand, Gemma and Mistral are still severely vulnerable to jailbreak templates. For ‘None’, CM improves AASR over the English set while CMP decreases it. However, both models react far more adversely to jailbreak templates, as shown by the very high AASR across all templates. Similar to ‘None’, AARR drops significantly with CMP for both models across the templates. Despite this compromise on language understanding with CMP, the AASR remains maintained across the templates, demonstrating the severe vulnerability of both models to classic red-teaming.
Takeaway for RQ.1: Multilingual safety alignment remains a major challenge, especially for models like Gemma and Mistral.
High AARR in Table 1 denotes that the LLMs, despite encountering inputs with nonsensical spellings, correctly interpret the prompts, while at the same time, being unable to trigger the safety filters effectively. Across all templates, ChatGPT and Llama maintain comparable AARR for all prompt sets, demonstrating that both models maintain comparable input understanding regardless of the prompt set. However, responses for CM and CMP are not as relevant for Gemma and Mistral, which report lower AARR than for English prompts. For the ‘None’ case, AASR for both models increases with CM; however, we see a slight drop in AARR for Gemma. The AARR drop in the English $\rightarrow$ CM transition is observed across the templates for Gemma. While Mistral nearly maintains and, in some cases, even improves AARR in this transition, both models undergo considerable drops in AARR in the CM $\rightarrow$ CMP transition across all configurations. These observations align with existing findings on multilingual proficiency, which indicate that ChatGPT outperforms Llama Zhou et al. (2024); Hendrycks and Dietterich (2019), and Llama in turn outperforms Gemma Thakur et al. (2024). Interestingly, we find in Table 1 that the best AARR scores are primarily all for ChatGPT, directly validating the above finding, and the overall best possible AASR jump from CM $\rightarrow$ CMP is also in the case of ChatGPT for the ‘None’ configuration, while maintaining almost perfect AARR in the transition. While the same argument for AASR is approximately true for Llama, AARR drops slightly, further validating the above finding.
Takeaway for RQ.2: The higher the multilingual proficiency of a model, the higher is the relevance of model responses and the more pronounced is the effect of our CMP strategy.
From the interpretability results in Figures 2– 4, we can infer that input tokenization plays an important role in determining whether the model generates a safe response or a harmful one. It is immediately evident that if the spelling or language of the sensitive word is altered, its tokens do not report high attribution scores for a safe response, thus resulting in harmful outputs.
Takeaway for RQ.3: Proper input tokenization could hold the key to more robust safety alignment design.
To test the generalization of our red-teaming approach, we also evaluate ChatGPT-4o-mini and Gemini-2.5-Flash-Image for the text-to-image generation task. The results in Table 2 show that for ChatGPT, both our red-teaming prompts and our novel jailbreak template, VisLM, are effective in jailbreaking the model and generating highly offensive, dangerous and harmful outputs. In fact, the AASR values are 25-65% higher for the image generation task than for the text generation task, across the prompt sets. While both models exhibit high AARR in all configurations, we observe a noticeable drop in AASR for Gemini in the ‘Base’ case with CM. Although the only case of a drop in AASR, other configurations with CM and CMP do not offer any major AASR jumps either. A more in-depth analysis reveals that for ‘Base’ with the English set, the model itself refuses to assist Yuan et al. (2024) with a large number of prompts. However, in both the English $\rightarrow$ CM and CM $\rightarrow$ CMP transitions, we notice that the number of refusals triggered by the model drops, but a larger number of generations are blocked by the API’s content moderation filter  [^3] See Appendix for more details. Google (2025b, a) due to their harmful nature. Interestingly, VisLM significantly reduces model refusals across all prompt-sets, indicating the effectiveness of our novel template as well as the vulnerability of both ChatGPT and Gemini to previously unseen jailbreak templates.
This generalisation to other output modalities shows how easy it is to jailbreak commercial black-box language models like ChatGPT and Gemini with simple modifications to input prompts. More importantly, since the outputs are of a visual nature, harm amplification is stronger Hao et al. (2024). We also note how certain categories of prompts, like “gore” are significantly easier to jailbreak as opposed to more niche ones, like “caste discrimination” which are relevant to a limited demographic.
Takeaway for RQ.4: Our red-teaming strategy generalizes to image generation with a higher AASR than for text generation.
Our findings from this study highlight three critical concerns for safety and alignment of LLMs.
Improved safety measures for all models: While Llama and ChatGPT show resistance to standard jailbreak attacks, Gemma and Mistral are easily jailbroken, even in English. This highlights the need for improved evaluation and alignment methods for these models, particularly given their widespread availability. In the multilingual and multimodal settings, especially with code-mixing, safety filters of all models degrade drastically. The success of our approach on Gemini-2.5-Flash-Image, a fairly recent and more advanced model, amplifies this issue even further, revealing the need for better efforts in multilingual multimodal safety and alignment.
Input tokenization determines safe outputs: If input tokens are out of vocabulary or perturbed, the safety filters get bypassed easily, but the model is able to interpret the instructions in the input prompt and generate harmful responses. Thus, a more robust tokenization strategy needs to be developed.
Template-based safety measures do not generalize well: The results from our experiments conducted across three open-source and two commercial models using code-mixing in the Hindi language show how template-based safety measures often fail and are not robust. Moreover, we show that even template-attack robustness can fail when faced with newer and stealthier templates. Therefore, we conclude that guardrails of models that are safety trained against template based attacks do not generalize well to attacks that deviate from set patterns, by including elements such as perturbations or code-mixing. This highlights the need for more generalizable safety training methods that are robust to deviations from standard attack patterns.

### v4

@@ 4 Results & Observations
We now present the results from our red-teaming experiments for all RQs described previously.
4.1 Success of CMP-RT (RQ1.)
@@
Table 1 reports the AASR for all prompt sets, models & jailbreak templates.
ChatGPT and Llama: The models are fairly robust to attacks in English, with AASR decreasing further when combined with the jailbreak templates. For ‘None’, AASR significantly increases in both the English $\rightarrow$ CM and CM $\rightarrow$ CMP transitions. While CMP substantially improves AASR for both models for the ‘None’ case, combining the CM or CMP prompts with the jailbreak templates again results in $\simeq$ 0 AASR in most cases, revealing strong alignment against template-based attacks.
Gemma and Mistral: Both models report very high AASR across the templates and prompt sets. For ‘None’ on the English set, both models already yield high AASR, showing severe vulnerability to classic template-based attacks despite evidence of harm. With the CM set for ‘None’, Gemma becomes highly complicit, while Mistral also shows a significant AASR jump. When combined with the templates, models reach up to 0.99 AASR for both the English and the CM sets. In the CM $\rightarrow$ CMP transition, AASR stays nearly the same across the templates, with a slight drop in the ‘None’ case.
We note that while ChatGPT and Llama are in general robust whereas Gemma and Mistral are brittle to template based-attacks, none of the templates show an obvious advantage over others across the configurations.
Significance Testing: Next, we evaluate the benefits of CM over English and CMP over CM using the Wilcoxon test Wilcoxon (1992) for each model and jailbreak template individually($p$-value = 0.050.05)2. We find that for CM is beneficial for None for Mistral, None and AIM for Llama and Gemma, and None, OM, Sandbox and AIM for ChatGPT. On the other hand, CMP provides benefits for ChatGPT and Llama in the None case, solidifying the resistance of ChatGPT and Llama against known template-based attacks.
We also establish the validity of the LLM-Judge, comparing its responses with those of human annotators. We obtain an ICC value of 0.87 for the ChatGPT English-set and 0.814 for the Gemma CMP-set responses, indicating a high degree of agreement Bartko (1966).
Defense Robustness: We find that while the OpenAI moderation API flags $\sim$61% of the harmful English prompts, it flags only 7.39% of the CMP prompts. Within each pair of safe & harmful prompt sets, we obtain comparable GPT-2 perplexity (PPL) values (Default: 57.30, Safe-Default: 73.7, English: 23.40, Safe-English: 26, CMP: 419.80 and Safe-CMP: 440.20). Comparable scores within a pair indicate how PPL fails to distinguish harmful inputs from benign ones. This implies that any PPL-based filtering would result in blocking benign and harmful alike, further validating the threat posed by CMP-RT, even against standard defenses.
Overall, these results establish CMP-RT as a realistic, non-trivial threat model, revealing under-generalization of existing alignment techniques and adversarial defenses. Next, we evaluate the input interpretability level of models with CMP-RT.
Metric Models Red-Teaming Strategy
Default-CSRT Hypothetical-CSRT CMP-RT (Ours)
AASR ChatGPT 0.10 0.23 0.50
Llama 0.15 0.31 0.63
Gemma 0.25 0.35 0.55
Mistral 0.40 0.52 0.68
AARR ChatGPT 0.44 0.56 0.99
Llama 0.24 0.43 0.95
Gemma 0.30 0.38 0.65
Mistral 0.22 0.34 0.74
Table 2: AASR and AARR scores for ChatGPT, Llama, Gemma, and Mistral for our Base and English prompt-sets converted to CSRT style inputs, against CMP-RT for the ‘None’ jailbreak-template. Metric-wise maximum values in each row are in bold.
4.2 Relevance of Generated Outputs (RQ2.)
@@
ChatGPT and Llama: Both models maintain very high AARR across all jailbreak templates, with AARR $\simeq$ 1 in many cases. Irrespective of the template, both models report AARR on CM and CMP sets comparable to the English set in most cases, denoting their ability to correctly interpret the prompts despite encountering inputs with nonsensical spellings, while being unable to trigger the safety filters at the same time. Thus, the majority of the responses are highly relevant to the prompt and, therefore, harmful.
Gemma and Mistral: Both models have generally high AARR scores on the English set across all templates. AARR drops significantly in the English $\rightarrow$ CM and CM $\rightarrow$ CMP transitions for Gemma and the CM $\rightarrow$ CMP transition for Mistral in most cases. Simultaneously, the AASR has a non-increasing trend except for ‘None’, demonstrating that despite compromised understanding with CMP, both models undergo only minor AASR drops despite, demonstrating the severe vulnerability of both models to classic red-teaming. For the ‘None’ case, AASR for both models increases with CM; however, we see a slight drop in AARR for Gemma. This English $\rightarrow$ CM AARR drop is observed across the templates in the case of Gemma. While Mistral nearly maintains and, in some cases, even improves AARR in this transition, both models undergo considerable drops in AARR in the CM $\rightarrow$ CMP transition across all configurations.
These observations align with existing findings on multilingual proficiency, which indicate that ChatGPT outperforms Llama Zhou et al. (2024); Hendrycks and Dietterich (2019), and Llama in turn outperforms Gemma Thakur et al. (2024). Interestingly, we find in Table 1 that the best AARR scores are primarily all for ChatGPT, directly validating the above finding, and the overall best possible AASR jump from CM $\rightarrow$ CMP is also in the case of ChatGPT for the ‘None’ configuration, while maintaining almost perfect AARR in the transition. While the same argument for AASR is approximately true for Llama, AARR drops slightly, further validating the above finding. This reveals that higher multilingual proficiency leads to more relevant response generation with CMP-RT.
4.2.1 Benchmarking CMP-RT
Table 3 shows that CMP-RT consistently outperforms both CSRT variants on AASR across all evaluated models, indicating stronger jailbreak effectiveness under identical conditions. More notably, CMP yields substantially higher AARR than both CSRT variants, demonstrating that CMP elicits harmful responses that remain far more aligned with the original prompt intent. We also note that across all models, AASR and AARR are strictly higher for Hypothetical-CSRT than Default-CSRT.
These findings demonstrate that CMP-RT preserves prompt intent significantly better than prior multilingual red-teaming strategies while eliciting harmful outputs, positioning it as a strong diagnostic probe for evaluating the generalization of safety alignment under realistic, non-canonical inputs.
4.3 Explaining Phonetic Perturbations (RQ3.)
@@ 4.2.1 Benchmarking CMP-RT
In Figures 24, we plot the sequence attribution scores using the method of Integrated Gradients Sundararajan et al. (2017), for a given input prompt in the English set, as well as its CM and CMP variants for the Llama-3-8B-Instruct LLM. We study the attribution scores for generating safe outputs– a higher token score implies higher contribution towards triggering the safety filter and vice versa. In Figure 2, we observe how the tokens of the prompt 2“hate”, “speech” and “discrimination” have a high attribution score at the embedding layer, 1st decoder layer and the 8th decoder layer. This means that these tokens are primarily responsible for generating the safe response from the LLM. “hate” and “speech” retain their high importance even at the 16th decoder layer. A similar observation can be made in Figure 3, for the same prompt written in a code-mixed (HI-EN) manner, where “hate” and “speech” are written in the English language. Thus, standard code-mixing may not be enough to bypass the safety filters of LLMs. On the other hand, in Figure 4, we note that the CMP variant generates radically different tokens– “hate” $\rightarrow$ “haet” tokenized as “ha” + “et” and, “discrimination” $\rightarrow$ “bhed bhav” tokenized as “b” + “hed” + “b” + “ha” + “av”. This leads to a change in the attribution scores as well, which are now low for the sensitive words and thus do not trigger the safety filters. A similar observation can be made for other prompts.
Thus, tokenization is exposed as a key safety vulnerability for transformers. More generally, this reveals a gap between pre-training and safety alignment. Pre-training on internet data likely exposes models to texts that include similar phonetically inspired perturbations as in CMP-RT, explaining the consistently high AARRs across our models. However, the high AASRs show how guardrails fail to activate despite excellent prompt interpretability, revealing that current safety training fails to incorporate such nuanced, real-world elements, despite advanced multilingual red-teaming and adversarial training strategies, leaving models exposed.
We conclude that phonetic perturbations lead to input tokenisation in a way that impacts the safety filters of LLMs, thus allowing attackers to generate harmful outputs.
4.4 CMP-RT for Image Generation (RQ4.)
@@
Table 3 shows the results of our image generation task. We present example generations from ChatGPT for each category of input prompt in Figure 5.
ChatGPT: The model is quite robust to English attacks in the Base (equivalent to ‘None’ in Table 1) case. AASR noticeably increases with CM, with a significant boost with CMP for both templates with consistently high AARR– similar to text generation. VisLM outperforms Base for all prompt sets.
Gemini-2.5-Flash-Image: For ‘Base’, AASR drops considerably in the English $\rightarrow$ CM transition, followed by a slight recovery with the CMP set, while ‘VisLM’ with CM and CMP offers only minor (although consistent) AASR jumps. An in-depth analysis reveals that for Base with the English set, the inherent model refusals Yuan et al. (2024) are much higher. However, for both the English $\rightarrow$ CM and CM $\rightarrow$ CMP transitions, the number of model refusals drops largely, but a larger number of harmful generations are blocked by the API’s moderation filter  [^2] See Appendix for more details. Google (2025b, a). Interestingly, VisLM significantly reduces model refusals across all prompt-sets, again achieving the highest AASR on the CMP set. AARRs remain consistently high across configurations.
Metric Models Jailbreak Templates
Base VisLM
@@
AASR ChatGPT 0.20 0.29 0.65 0.35 0.45 0.78
Gemini-2.5 0.30 0.19 0.25 0.38 0.40 0.43
Nano Banana Pro 0.45 0.48 0.65 0.63 0.69 0.76
AARR ChatGPT 0.93 0.98 0.95 1.00 0.98 0.94
Gemini-2.5 0.97 0.96 0.94 0.98 0.98 0.96
Nano Banana Pro 0.94 0.97 0.93 0.96 0.89 0.91
@@
Table 3: AASR and AARR scores for ChatGPT, Gemini-2.5-Flash-Image (API), and Nano Banana Pro across Base and VisLM jailbreak templates and all input sets (Eng, CM, CMP). Metric-wise maximum values in each column are in bold.
Nano Banana Pro: CMP-RT effectively probes Nano Banana Pro. In contrast to Gemini-2.5, both CM and CMP consistently increase AASR across the templates, with the highest scores achieved yet again with VisLM on the CMP set. AARRs remain consistently high across configurations.
For all three models, VisLM achieves substantially higher AASRs over Base, contradicting our text generation finding where advanced models demonstrate resistance to template-based attacks. This shows that while system level guardrails protect models from existing template-based attacks, newer, stealthier templates still reveal under-generalization and scope for stronger control..
Table 4 presents the AASR per input prompt category for the image generation task. We note the high gore generation tendencies of ChatGPT and Nano Banana. For Gemini-2.5, gore generation suppression are likely higher at the platform level, making SM toxicity its best category overall (which is also the 2nd best Nano Banana category).
Models Category Base VisLM
Eng CM CMP Eng CM CMP
ChatGPT Religious Hate 0.10 0.25 0.75 0.30 0.40 0.90
Gore 0.45 0.70 0.85 0.75 0.70 0.90
Self-Harm 0.15 0.20 0.85 0.20 0.45 0.95
Casteist Hate 0.10 0.10 0.30 0.15 0.20 0.25
SM Toxicity 0.30 0.35 0.60 0.50 0.60 0.80
Gemini-2.5 Religious Hate 0.40 0.15 0.30 0.55 0.50 0.65
Gore 0.10 0.15 0.15 0.05 0.05 0.15
Self-Harm 0.40 0.25 0.20 0.50 0.50 0.45
Casteist Hate 0.25 0.10 0.15 0.30 0.45 0.45
SM Toxicity 0.45 0.40 0.50 0.65 0.65 0.60
Nano Banana Pro Religious Hate 0.35 0.30 0.55 0.40 0.65 0.60
Gore 0.80 0.70 0.90 0.90 0.75 0.90
Self-Harm 0.50 0.70 0.65 0.60 0.80 0.70
Casteist Hate 0.05 0.10 0.45 0.35 0.45 0.50
SM Toxicity 0.45 0.75 0.80 0.85 0.75 0.85
Table 4: Category-wise AASR for the image generation task for ChatGPT, Gemini-2.5-Flash-Image (API) and Nano Banana Pro across both jailbreak templates and all input sets, i.e., Standard English prompts, CM prompts and CMP prompts. Template-wise maximum values for each model-prompt set pair are in bold.
Our fine-tuning based CMP-RT automation experiment results in AASR and AARR of 0.46 and 0.77 on the 521 AdvBench examples evaluated on Llama-3-8B-Instruct. A qualitative inspection indicates that prompt transformations up to the Default $\rightarrow$ English $\rightarrow$ CM stages generally preserve the intended semantics. However, the final CM $\rightarrow$ CMP conversion can introduce over-corruption of sensitive tokens, and at times even random corruptions that may not reflect the true nature of a phonetic style attack. This reflects in the AASR and AARR reduction relative to the manual baseline. Nevertheless, the automatically generated CMP prompts retain non-trivial jailbreak success and relevance, establishing CMP as scalable threat model.
This generalization of CMP-RT across modalities, SOTA models like ChatGPT and Gemini-3-Pro and scalability through simple fine-tuning reveals the severity of tokenization as a fundamental vulnerability and a real-world exploit. Using our hand-crafted, high-quality CMP dataset as seed, we demonstrate how a simple tokenizer exploit exposes models to attacks at much larger scale.